OpenStudy (pottersheep):

I really don't know how to do this quadratic equation question :( Please help

OpenStudy (anonymous):

?

OpenStudy (pottersheep):

The height (h) in metres of a football at time "t" is given by h = 23t - 5t^2. Does the football ever reach 27m ?

OpenStudy (anonymous):

so your job is it solve \[27=23t-5t^2\] if you can

OpenStudy (pottersheep):

yup

OpenStudy (anonymous):

first step is to write as \[5t^2-23t+27=0\] and now since it doesn't even ask for a solution you can just find \[b^2-4ac\] and see if it is positive or negative

OpenStudy (pottersheep):

ohhhhhhhhhhhhhhh and if it's negative then the answer is "no" right?

OpenStudy (anonymous):

in this case you get \[23^2-4\times 5\times 27=-11\] so there is no solution

OpenStudy (pottersheep):

I mean if the radical has a negative number in in..

OpenStudy (anonymous):

right

OpenStudy (pottersheep):

Ahhh okay thank you thank you :))))

OpenStudy (anonymous):

you could also just solve \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] but you will see you get a negative number under the radical, so no answer

OpenStudy (pottersheep):

Right ! I totally get it ~ thanks so much :D:

OpenStudy (anonymous):

yw